At a turnover of 200 million MU. and a profit rate calculated on turnover of 15%, if the rate of profit becomes 7.5% while maintaining the same turnover, determine how the profit and the cost of production change.


1
Expert's answer
2022-02-14T15:26:57-0500

If the turnover is 200 million,

Profit Of 15%"=\frac{15}{100}×200000000=30000000"

Cost"=Turnover-Total Profit=200000000-30000000=170000000"

Profit Of 7.5%"=\frac{7.5}{100}×200000000=15000000"

New cost"=200000000-15000000=185000000"

Therefore the profit change "=30000000-15000000=15000000"

Cost is changing by

"185000000-170000000=15000000"